OpenStudy (dls):

x^3+y^3+3axy=0 dy/dx=?

OpenStudy (dls):

\[\LARGE 3x^{2}+3y^{2} \frac{dy}{dx}-3ax=0\] this gives..finally.. \[\LARGE \frac{dy}{dx}=\frac{x(a-x)}{y^{2}} \] I doubt this because im quite unsure about the derivative of 3axy,we are differentiating wrt x or y? if its x then 3ay(dx//dx) not sure..

OpenStudy (dls):

@hartnn

hartnn (hartnn):

you'll need product rule for 3axy

OpenStudy (dls):

I thought product rule was used for function times function

OpenStudy (dls):

thats a variable,isn't it

hartnn (hartnn):

y is the function of x, x is the function of x.

OpenStudy (dls):

:O

hartnn (hartnn):

(xy)' = x'y+xy' = y+xy'

OpenStudy (dls):

so how will we find for 3axy?

hartnn (hartnn):

3a (y+xy')

OpenStudy (dls):

\[\LARGE 3ay(\frac{dx}{dx}) |||| 3ax(\frac {dy}{dx}) \]

hartnn (hartnn):

dx/dx=1

OpenStudy (dls):

yeah!

hartnn (hartnn):

and i assume ||| is +

OpenStudy (dls):

yes

OpenStudy (dls):

thats it?:O

hartnn (hartnn):

so, dy/dx will change..

OpenStudy (dls):

what do u mean?

hartnn (hartnn):

i mean recalculate dy/dx

OpenStudy (dls):

\[\LARGE \frac{dy}{dx}= \frac{-3(ay+x^{2})}{3(y^{2}-ax)} \]

hartnn (hartnn):

-ax ?

OpenStudy (dls):

sorry + printing mistake :P

hartnn (hartnn):

then ok.

OpenStudy (dls):

:D